OpenStudy (anonymous):

Help @ModernPraise97 . Here it is There are 20 kilograms of a solution that is 10% salt. How many kilograms of salt must be added to make a solution that is 25% salt?

OpenStudy (beccaboo333):

I can say it in a simple way.. half of 20 is 10 which is 5%. 10% is 20 20+20+10= 30

OpenStudy (beccaboo333):

I'm not very good at explaining things. Sorry :(

OpenStudy (anonymous):

Keep going Adam beccaboo333 did not help me.:)

OpenStudy (beccaboo333):

I gave you the answer lol.. Sorry I just am not good at explaining math.

OpenStudy (anonymous):

how much is 10% of 20 kilograms? 2 Kilograms. There for we know that 5% is equal to 1 kilogram. 10% = 2 K 5% = 1 K to get to 25% we need to add the 10 and the 5.. sooo we need to add the 2K and the 1K which would be 3. There for... 10% solution has 2 K 25% solution has 3 K how many K do we need to add to get from 10% to 25%

OpenStudy (anonymous):

sorry it took so long lol

OpenStudy (anonymous):

That is ok beccaboo333.:)

OpenStudy (anonymous):

In answer to your question Adam- I am not really sure. Give me one sec...lol

OpenStudy (anonymous):

to get from 10% to 25% we need 15% To get from 2 to 3 we need x

OpenStudy (anonymous):

Ohhhh I see. Keep going...:)

OpenStudy (anonymous):

3 - 2 = x

OpenStudy (anonymous):

Hey Adam, I see what u are doing, but I am a little confused.:)

OpenStudy (anonymous):

well I figure that 2 Kilograms of salt is 10% of the 20 Kilogram solution... and 3 Kilograms of salt is 25% of the 20 kilogram solution right?

OpenStudy (anonymous):

EXCUSE ME! wow fail of the day... 5 KILOGRAMS is 25%.... -facepalm-

OpenStudy (anonymous):

Your fine...lol Everyone makes a mistake:)

OpenStudy (anonymous):

Let me see if I can work this...

OpenStudy (anonymous):

In answer to your original question- How many K do we need to add to get from 10% to 25%... 7 K ?

OpenStudy (anonymous):

to get from 2 K to 5 K we need 3 K...

OpenStudy (anonymous):

Ha Ha Ha I'm sorry not thinking:)

OpenStudy (anonymous):

its all good :)

OpenStudy (anonymous):

I'm thinking...lol:)

OpenStudy (anonymous):

take your time :)

OpenStudy (anonymous):

I was thinking that i could write the formula out to be: 0.10(20)=0.25(20+x) Seeing as I have changed the percents to decimal fractions in the past. What do you think???

OpenStudy (anonymous):

sounds good to me

OpenStudy (anonymous):

Well I guess I will work it.:) Give me one minute...:)

OpenStudy (anonymous):

That totally doesn't work. Do you have an suggestions??? This is what happened: 0.10(20)=0.25(20+x) 2=5+x x=-3 Could you help me right out a problem in this form though?

OpenStudy (anonymous):

it would be .10(20) = .25(20 - x)

OpenStudy (anonymous):

OK so... .10(20)=.25(20-x) 2=5-0.25x -3=-0.25x x=12 Is this correct???

OpenStudy (anonymous):

hmmm.. there is something up with my solving this problem... I mean I have the answer I just can't set it up like you want me too...

OpenStudy (anonymous):

So the answer isn't 12???

OpenStudy (anonymous):

no...

OpenStudy (anonymous):

Hey Adam, I need to eat dinner so if you reply I'm not gone. I will send a text your way when I'm finished. I will think about the problem and be back in no time.Thanks again for your help!!! Snow:)

OpenStudy (anonymous):

no prob :)